OpenStudy (anonymous):

Ian solved the system of equations by making a substitution for y in the second equation. What equation did he get as a result of that substitution? y = –3x + 1 5x – 2y = 31

OpenStudy (anonymous):

5x + 6x + 1 = 31 5x + 6x – 2 = 31 –15x + 5x = 31 –15x – x = 31

OpenStudy (owlcoffee):

since y is already solved for in the first equation, we can take the second equation and replace that value of "y" on it, so it would look like this; \[5x-2(-3x+1)=31\] and doing the distributive: \[5x+6x-2=31\] and operating similar terms: \[11x-33=0\] And that counts as the equation he would obtain if he did obtain as a result of the sustitution.